OpenStudy (i0iz0gangster):

Can someone show me how 2 simplify (2 √5 +3 √7)² step by step

OpenStudy (anonymous):

You'll want to FOIL this out. But first, rewrite it like this: \[(2\sqrt{5} + 3\sqrt{7})(2\sqrt{5} + 3\sqrt{7})\]Now, you can FOIL this out as you normally would.

OpenStudy (anonymous):

Apply (a + b)^2 formula to it

OpenStudy (i0iz0gangster):

so it would 2 sqrt 18 sqrt 7?

OpenStudy (anonymous):

It would be 103 + 12underroot 35

OpenStudy (i0iz0gangster):

do you mind explaining I still cant figure out how you got that answer...

OpenStudy (anonymous):

The formula goes a^2 + 2ab + b^2 Hence place the first term as a Second term as b Plug the value get the answer. Simple

OpenStudy (lncognlto):

2sr5 + 3sr7 ^2 =2sr5(2sr5) + 2sr5(3sr7) + 3sr7(3sr7) =4(5) + 6sr35 + 9(7) =83 + 6 sr35

OpenStudy (i0iz0gangster):

83+6 sr 35 is the final answer or should I complete it ?

OpenStudy (anonymous):

Great u almost got it but u forgot to multiply the middle term by 2

OpenStudy (anonymous):

My bad its 83 ur right

OpenStudy (anonymous):

The final is 83 +12sr35

OpenStudy (i0iz0gangster):

thank you all for the help god bless :)